OpenStudy (anonymous):

In a power series, what is the relationship between the x and the radius of convergence? For example, if the radius of convergence of Cn(x^n) is R, then what is the relationship that makes the radius of convergence for Cn((x/2)^n) equal to 2R? Does dividing the "x" term in the series by a number such as 3 make the radius of convergence 3 times larger?

OpenStudy (amistre64):

in essense: |f(x)| = L such that R = inverse of f(x) i believe

OpenStudy (amistre64):

\[\lim~(2x)^n\] \[\lim~2x\] \[|2x|\lim~1\] \[|2x|<1\] \[|x|<\frac12~:~R=\frac12\]

OpenStudy (anonymous):

use ratio test to determin convergence: Notice that the convergence is taken at \(x_0=0\) \(\frac{C_{n+1}x^{n+1}}{C_nx^n}=\frac{C_{n+1}}{C_n}x\) series will converge if \(\frac{C_{n+1}}{C_n}x<1\) or \(x<\frac{C_{n}}{C_{n+1}}\)